The length of a rectangular playing field is 5 ft less than twice its width. If the perimeter of the playing field is 230 ft, find the length and width of the field.
Answer by pvaka(23)

P=2L+2W
L=x,W=y
x=2y-5, P=230
--------------------------------------
2x+2y=230,x=2y-5>>(Substitute 2y-5 for x)
2(2y-5)+2y=230
4y-10+2y=230
6y-10+10=230+10
6y=240
6y/6=240/6
y=40
x=2y-5
x=2(40)-5
x=80-5
x=75
Length=75 feet,width=40 ft